The Length of a rectangle is 4 less than 3 times the width. if the perimeter is 40, find the dimensions of the rectangle
Answer by addingup(3677)

L = 3W-4
and
2L+2W = 40 Substitute for L:
2(3W-4)+2W = 40
6W-8+2W = 40
8W = 48
W = 6 This is the width, and the Length:
L = 3(6)-4 = 14
```````````````````````````````````````````````
Check:
2(6)+2(14)= 40
12+28 = 40 Correct.